OpenStudy (anonymous):

Let f(x) = -2x - 7 and g(x) = -4x + 3. Find (f o g)(-5)

jimthompson5910 (jim_thompson5910):

(f o g)(-5) is the same as f(g(-5)) the first task is to find g(-5). Do you know how to do this?

OpenStudy (anonymous):

A. -53 C.23 D.3

OpenStudy (anonymous):

No i have no clue

OpenStudy (anonymous):

@satellite73

OpenStudy (anonymous):

Could you help @jim_thompson5910

OpenStudy (anonymous):

@SolomonZelman

jimthompson5910 (jim_thompson5910):

g(x) = -4x + 3 g(-5) = -4(-5) + 3 ... replace every x with -5 g(-5) = ??

OpenStudy (misssunshinexxoxo):

On OS we help people learn; Jim is trying to make sure you are understanding this

OpenStudy (anonymous):

Oh i know and i defiantly want the help

OpenStudy (usukidoll):

(f o g)(-5) well first let's do (f o g ) which means that the g(x) equation is placed inside the f(x) equation

OpenStudy (anonymous):

20+3

OpenStudy (usukidoll):

whoa... let's do this step by step

OpenStudy (anonymous):

Oh sorry i was replying to what @jim_thompson5910 started

jimthompson5910 (jim_thompson5910):

so you should find that g(-5) = 23 agreed?

OpenStudy (anonymous):

correct

jimthompson5910 (jim_thompson5910):

f(g(-5)) becomes f(23) after you replace g(-5) with 23

jimthompson5910 (jim_thompson5910):

now follow the same basic steps to calculate f(23)

OpenStudy (usukidoll):

ok it's ok didn't see that :) anyway given the functions f(x) = -2x-7 and g(x) = -4x+3 we need to place g(x) =-4x+3 into the f(x) so for every x in f(x) the g(x) takes it's place

OpenStudy (usukidoll):

so the end result should be -2(insert g(x) function here)-7

OpenStudy (usukidoll):

then combine like terms and evaluate when x = -5

OpenStudy (anonymous):

im lost

OpenStudy (anonymous):

Is the answer 23?

jimthompson5910 (jim_thompson5910):

no

OpenStudy (usukidoll):

I agree with jim there's more to be done...

OpenStudy (anonymous):

ok im ready to learn the rest

OpenStudy (usukidoll):

k.. so did you already place your g(x) into your entire f(x) function?

OpenStudy (anonymous):

No i got lost after we got f(23)

jimthompson5910 (jim_thompson5910):

To calculate f(23), replace every x in f(x) with 23 f(x) = -2x - 7 f(23) = -2*23 - 7 f(23) = ??

OpenStudy (usukidoll):

that means x = 23 evaluate the f(x) function

jimthompson5910 (jim_thompson5910):

it's just like the g(x) one

OpenStudy (anonymous):

-46

OpenStudy (anonymous):

f(23)=-46-7

jimthompson5910 (jim_thompson5910):

keep going

OpenStudy (usukidoll):

almost there

OpenStudy (anonymous):

46-7=39

OpenStudy (usukidoll):

we have two negative numbers so the negative numbers are supposed to increase, not decrease

OpenStudy (anonymous):

opps wrong

OpenStudy (anonymous):

-53

OpenStudy (usukidoll):

yes!

jimthompson5910 (jim_thompson5910):

good

OpenStudy (usukidoll):

do you want to see another method of solving this equation as a reference?

OpenStudy (anonymous):

yay we did it

OpenStudy (anonymous):

yes that would be great

OpenStudy (usukidoll):

ok so here's how I would tackle this problem (f o g ) (-5) first I will tackle f o g portion and then evaluate when x = -5 so f o g means that we place our g(x) function inside all x's of the f(x) function f(x) = -2x-7 and g(x) = -4x+3 f o g -2(-4x+3)-7 distribute the -2 8x-6-7 8x-13 so our f o g = 8x-13 now evaluating when x = -5 f o g (-5) = 8(-5)-13 = -40-13 = -53

OpenStudy (anonymous):

Can you guys help me solve another @UsukiDoll @jim_thompson5910

OpenStudy (usukidoll):

sure

OpenStudy (anonymous):

Let f(x) = x + 2 and g(x) = x2 Find ( g o f)(-5)

OpenStudy (usukidoll):

\[\large f(x) = x+2, g(x) = x^2 \]

OpenStudy (usukidoll):

this is similar to the previous problem. while our x still remains the same g o f means that we place our entire f(x) function inside the x's of the g(x) function so g o f = (insert f(x) here)^2

OpenStudy (anonymous):

so is x -2?

OpenStudy (usukidoll):

our f(x) = x+2 and we need place it inside g(x). Since g(x) =x^2 g o f = (f(x) function)^2

OpenStudy (anonymous):

so X^2+X+2?

OpenStudy (usukidoll):

\[g o f =(x+2)^2 \] that second power means write (x+2) two times after that, we need to use FOIL to expand

OpenStudy (anonymous):

your gonna have to hold my hand on this one cuz im totally lost lol im sorry

OpenStudy (usukidoll):

alirght.. so we write (x+2) two times (x+2)^2 means (x+2)(x+2) now using FOIL (x)(x)+2(x)+2(x)+(2)(2) so what's 2 x 2 2x + 2x and x times x I guess that's the closest way of notating multiplication

OpenStudy (anonymous):

Ok so where do we go from this?

OpenStudy (usukidoll):

first tell me what 2 x 2 is then tell me what 2x+2x is .. I will factor out the x to make it easier x(2+2)

OpenStudy (anonymous):

2x2=4

OpenStudy (anonymous):

2x+2x=4X

OpenStudy (usukidoll):

mhm last one what is (x)(x) . I'll make this easier by rewriting it and using the addition exponential rule, \[\large (x)(x) \rightarrow x^1x^1 \rightarrow x^{1+1}\] so you just have to solve for the exponent part. what is 1+ 1

OpenStudy (anonymous):

2

OpenStudy (anonymous):

\[x ^{2}\]

OpenStudy (usukidoll):

yes so now we have \[g o f =x^2+4x+4 \]

OpenStudy (usukidoll):

so now we evaluate this when x = -5. so replace all x's with -5

OpenStudy (anonymous):

25-20+4=9

OpenStudy (usukidoll):

\[g o f (-5) =(-5)^2+4(-5)+4\] \[g o f (-5) =25-20+4 \rightarrow 5+4 = 9 \] correct

OpenStudy (anonymous):

Thank you very much